OpenStudy (anonymous):

What is the best way to classify each equation? Column A Column B 1. 22 - 3x + 7x = 4(x + 5) 2.6(2x - 3) = 3(3x - 5) 3.6x - 14 = 2(3x - 7) 4.6x + 3 (x - 4) = 8(x - 3) A. identity B. contradiction C. neither

OpenStudy (texaschic101):

identity equations are always true no matter what you sub in for the variable...so basically, identity equations have infinite solutions. contradiction equations have no solution. so lets get to work.... 22 - 3x + 7x = 4(x + 5) -- distribute through the parenthesis 22 - 3x + 7x = 4x + 20 -- simplify 22 + 4x = 4x + 20 -- subtract 4x from both sides 22 = 20 --- incorrect....no solution...contradictory 6(2x - 3) = 3(3x - 5) -- distribute through the parenthesis 12x - 18 = 9x - 15 -- subtract 9x from both sides 12x - 9x - 18 = -15 -- add 18 to both sides 12x - 9x = -15 + 18 3x = 3 x = 1.....this tells me that there is 1 solution....so it is neither 6x - 14 = 2(3x - 7) -- distribute through the parenthesis 6x - 14 = 6x - 14 0 = 0...correct.....this means there is infinite solutions.... identity 6x + 3(x - 4) = 8(x - 3) -- distribute through the parenthesis 6x + 3x - 12 = 8x - 24 9x - 12 = 8x - 24 --- subtract 8x from both sides 9x - 8x - 12 = -24 -- add 12 to both sides 9x - 8x = -24 + 12 x = - 12....this tells me there is 1 solution...so this one is neither also